7. A 1300 kg car drives up a 17 m high hill (the elevation of the hill is 17 meters). During the drive two nonconservative forces do work on the car: the force of friction and the force generated by the car’s engine. The work done by friction adds 331 kJ to the internal energy. The work done by the engine is 634 kJ.

(a) What is the change in the car’s kinetic energy?

(b) If the car started up the hill with a speed of 20 m/s, what is the car’s speed at the top of the hill?

8. A 1.80 kg block slides towards a spring on a rough horizontal surface with coefficient of friction µk = 0.56. As the block reaches the spring it is traveling at 2.0 m/s. The block continues to move until the spring brings it to a stop. In the process the spring compresses 11 cm. What is the force constant k of the spring?
Answer #1

7

a) change in kE = 634 - 331 - mg h = 303 - 1.3* 9.8* 17 = 86.42 kJ

b)

0.5 m ( v^2 - u^2) = change in kE

0.5* 1300* ( v^2 - 20^2) = 86.42* 10^3

v = 23.09 m/s

========

plz post other problem separately

Comment before rate in case any doubt, will reply for sure.. goodluck